OpenStudy (anonymous):

A television network is about to telecast a new television show. Before a show is launched, the network airs a pilot episode and receives a report assessing favorable or unfavorable viewer response. In the past, 60% of the network's shows have received a favorable response from viewers, and 40% have received an unfavorable response. If 50% of the network’s shows have received a favorable response and have been successful, and 30% of the network’s shows have received an unfavorable response and have been successful, what is the probability that this new show will be successful if it receives a

OpenStudy (anonymous):

favorable response.

OpenStudy (anonymous):

(A).41, (B).53, (C).67, (D).70, (E).83

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Can i ask how?

OpenStudy (vivek3461):

If you take 100 shows 60 shows got favorable response and in that 50 shows were successful. So probability for a show to be successful if it got a favorable response is = 50/60 = 0.83

OpenStudy (anonymous):

I'll take your word for it. Thanks!!

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@Vivek3461 would you be able to help with another?

OpenStudy (vivek3461):

Yes

OpenStudy (anonymous):

Of all the Sunny Club members in a particular city, 25% prefer swimming on weekends and 75% prefer swimming on weekdays. 10% of the members in the city prefer swimming on weekends and are female. 55% of the members in the city prefer swimming on weekdays and are female. What is the probability that a randomly selected club member is female, given that the person prefers swimming on weekends?

OpenStudy (anonymous):

(A).19, (B).20, (C).24, (D).40, (E).55

OpenStudy (vivek3461):

Probability is = 10 / 25 ( no of female who likes to swim on weekends / Total no of persons like to swim on weekends )